OpenStudy (anonymous):

Giuliana has 22 quarters and dollar coins worth a total of $10.75. Which equation can be used to find d, the number of dollar coins Giuliana has? A. d – 22 + 0.25d = 10.75 B. 0.25d + 22 – d = 10.75 C. 0.25(22 – d) + d = 10.75 D. d + 0.25(d – 22) = 10.75

Vocaloid (vocaloid):

hint: d = number of dollar coins q = number of quarters d + q = 22, since there are 22 total coins we can rewrite this as q = 22-d d + 0.25q = 10.75, which is the total value of the coins substitute "22-d" for q